# Wireless Hearables With Programmable Speech AI Accelerators

### v1

@@ 1. Introduction
  • Tiny, low-power AI accelerator hardware (syn, 2025; gre, 2023) is inherently far more constrained than GPUs and general-purpose embedded CPUs like the Raspberry Pi. On the other hand, effective neural architectures for speech enhancement rely on components such as convolutions (Luo and Mesgarani, 2019b), LSTMs (Luo et al., 2020), transformers (Subakan et al., 2021a), and state-space layers (Li and Chen, 2024). These models prioritize speech quality over real-time, on-device, or low-power constraints — indeed, transformers and state-space models exceed our target hardware’s runtime and memory limits. While efforts like TinyLSTM (Fedorov et al., 2020) and its successor TinyDenoiser (Rusci et al., 2022) aim for on-device deployment, their 25 ms algorithmic latency falls short of the low-latency requirements for enhanced hearing applications.
@@
  • Real-time on-device efficient neural network. Prior on-device models like TinyDenoiser (Rusci et al., 2022) are constrained by their limited complexity and inability to achieve effective performance with sub-10ms latencies. Instead, we begin with dual-path models that process audio in the time-frequency (TF) domain, treating time and frequency components as sequences using recurrent networks (Yang et al., 2024; Chao et al., 2024). These architectures achieve state-of-the-art performance by modeling both temporal and frequency relationships. However, they are computationally expensive, as running recurrent networks across each frequency prevents real-time operation. Thus, they cannot run in real time on our hardware. To address this, our proposed architecture compresses frequencies using a linear layer and replaces LSTMs with Gated Recurrent Units (GRUs). These architectural changes and others listed in §2.2.2 achieve real-time performance by reducing runtime while outperforming TinyDenoiser for the speech denoising task.
  • Hardware-software co-design. Floating-point speech denoising models cannot achieve real-time operation on our target accelerator hardware. To address this, we explore various quantized model configurations and use quantization-aware training (QAT) to simulate non-linear quantization errors and fine-tune the network to mitigate error propagation across layers. This further narrows the performance gap between our mixed-precision and fully floating-point speech models.
@@ 2.1.1. Stacked printed circuits boards
  • PERIPH: A 4-layer rigid board contains peripheral devices, including an AP Memory APS256XXN 256 Mbit RAM chip and an SSM6515 ultra-low-power I2S DAC and amplifier. The AI accelerator can control power to the RAM chip using a load switch (FPF1204UCX) when the RAM is not needed. Additionally, this board includes a NOR flash (Macronix MX25UW12845G) and an IMU (BMI323), which are reserved for future research. A right-angled 1 mm pitch connector is used to connect a Phonak Audeo Marvel M Receiver-in-canal (RIC), allowing audio playback into the ear canal. The board connects to the MIC circuit using a Molex 5050661022 connector.
@@ 2.2. Efficient Streaming Neural Network
As shown in Fig. 3A, end-to-end latency is defined as the time for a single audio sample to travel from the microphone input buffer, through the speech enhancement network, and into the speaker output buffer. This latency consists of two components: algorithmic latency, introduced by factors such as chunk size, lookahead, and overlap-add; and hardware latency, which accounts for the computation and I/O time required to process each chunk (Wang et al., 2022a).
@@ 2.2.2. Neural Network Architecture
Dual-window approach for low-algorithmic latency. As shown in Fig. 3A, to achieve low algorithmic latency, we process the input audio $x_{i}$ in small chunks of size $L_{C}$ (ms), with an additional lookahead of $L_{F}$ (ms) and a lookback of $L_{B}$ (ms); we set $L_{C}=6ms$, $L_{B}=6ms$ and $L_{F}=4ms$. Lookahead and lookback padding improve frequency resolution, which is important for enhancement performance (Wang et al., 2022a). However, increasing the lookahead and lookback padding also introduce higher algorithmic latency.
@@
To eliminate the algorithmic latency introduced by lookback padding, we adopt a dual-window-size approach (Wang et al., 2022a; Mauler and Martin, 2007), where the synthesis window is shorter than the analysis window, as shown in Fig. 3D. By discarding the lookback samples after the inverse DFT and before applying the synthesis window, the overlap-add process for the current chunk no longer depends on the next chunk’s output. We use a rectangular window for analysis and use a specific function for the synthesis window $s$, described below, valid for $L_{C}>L_{F}$, to achieve perfect signal reconstruction (Griffin and Lim, 1984):
@@
Cache state management. When processing a stream of continuous chunks, many values can be reused to avoid redundant computations. To achieve this, we maintain and update a cache state $h_{i}$ at every inference step. As shown in Fig. 3C, our cache state $h_{i}$ consists of four components: 1) Since the kernel of the first 2D causal convolution layer requires prior chunks, we can avoid recomputing the STFT frames of these prior chunks by caching and reusing them, as illustrated in the figure. 2) Similarly, we store the output of the sequence of dual-path modules from previous chunks, allowing us to compute the causal 2D deconvolution more efficiently. 3) Since computing the ISTFT for the current chunk depends on previous chunks, we maintain a buffer for the intermediate outputs of the 2D deconvolution layer. 4) We store the hidden and cell states of the temporal unidirectional LSTM for each dual-path module. This allows us to fully leverage the long-term receptive field of the recurrent network.
@@ 2.2.3. Training for Real-world Generalization
In practice, each device runs its own network, processing audio independently. Thus, we train two separate networks -— one for the left channel and one for the right.
Fine-tuning with motion and colored noise. In the previous stage, the sound source and wearer’s head were assumed to be static. However, real-world scenarios involve motion. To address this, we fine-tune the model with time-varying $\theta$ and $\phi$, following the motion simulation method from  (Veluri et al., 2024).
Specifically, the source’s position updates every 25 ms, with a 2.5% probability of triggering a random motion event. When triggered, the angular velocity in the azimuthal and polar directions is sampled from [$\pi/6,\pi/2$] rad/s, and the source moves at this velocity for a random duration between 0.1 and 1 s, during which no other motion event is triggered. The Steam Audio SDK (SDK, 2023) is used for motion trajectory simulation, creating a mix of stationary and moving segments within the same audio clip. Since BRIR datasets contain recordings at discrete positions, we approximate intermediate positions using nearest-neighbor BRIR selection based on the current $[\theta(t),\phi(t)]$.
We also augment the mixture signal with white, pink, and brown noise, mimicking real-world noise sources like microphone thermal noise and HVAC systems. White noise is generated with a standard deviation sampled from [0,0.002), while pink and brown noise are created using the Python colorednoise library and scaled by a random factor from [0,0.05]. We fine-tune the model for 100 additional epochs, incorporating motion simulation and noise augmentation. Instead of SNR loss, we use a multi-resolution spectrogram loss (Yamamoto et al., 2020; Steinmetz and Reiss, 2020). Optimization uses AdamW with a gradient clipping of 0.1, and a ReduceLROnPlateau scheduler (patience = 5, factor = 0.5) starting at 1e-3 learning rate.
@@ 2.3. Hardware-Software Co-Design
Efficient DSP and neural network inference rely on GAP9’s compute cluster —- 9 RISC-V cores and the NE16 accelerator. The NE16 is optimized for streamed multiply-accumulate operations but only supports up to 8-bit quantized weights and 8-/16-bit activations. The RISC-V cores handle both integer and floating-point operations but are less efficient for neural networks. Thus, the network must be carefully designed by identifying the parts of the network that can be efficiently run on the NE16 with quantization without degrading the output speech quality.
@@
[…]
@@
A key aspect of uniform quantization is choosing the scaling factor $S$. It defines how a given range $[\alpha,\beta]$ of real values $r$ is divided into discrete partitions (Gholami et al., 2022): $S=\frac{\beta-\alpha}{2^{b}-1}$. In our implementation, we use the Min-Max Moving Average Observer (Cohen et al., 2023) to determine $[\alpha,\beta]$.
In asymmetric quantization, the range $[\alpha,\beta]$ is not necessarily symmetric with respect to the origin, i.e., $-\alpha\neq\beta$. In symmetric quantization, in contrast, a symmetric clipping range ($-\alpha=\beta$) is used and its scaling factor $S$ is computed as, $S=\frac{\text{max}(-\alpha,\beta)}{2^{b-1}-1}$ (Gholami et al., 2022), with the zero-point set to $Z=0$.
@@ 2.3.3. Mixed-Precision Configuration
One common method for quantizing a trained floating-point neural network is Post-Training Quantization (PTQ). This involves: 1) calibrating the quantization range using a representative dataset, and 2) quantizing all tensors in the network to a lower-precision format.
@@ 2.3.4. Quantization-Aware Training for Better Performance
Fine-tuning: The mixed-precision model is fine-tuned for 30 epochs using the FQSE framework (Cohen et al., 2023).
@@ 3.1. Benchmark results
Applying mixed-precision quantization helps recover 0.36 dB of SISDRi for TinyDenoiser and 2.6 dB for our model. However, the performance loss compared to the floating-point model remains non-negligible. After QAT, TinyDenoiser achieves an SISDRi of 5.97 dB, while our model reaches 8.19 dB, demonstrating the effectiveness of QAT, especially for deeper and more complex networks. Our mixed-precision model with QAT achieves the best performance within hardware constraints. Considering runtime, memory, and power consumption, our model balances performance and memory efficiency while maintaining real-time requirements ($<6~{}ms$) and power constraints ($71.6~{}mW$), trading some runtime and power efficiency for improved speech enhancement.
@@
In Fig. 4, we plot the output SISDR as a function of input SISDR for both the floating-point model and the mixed-precision QAT model. The results show that QAT reduces the performance gap between our floating-point and quantized models across all input noise levels. An interesting observation is that the performance gap increases with higher input SISDR. As input SISDR increases, the output speech quality becomes clearer, making quantization noise more dominant and its effects more pronounced compared to scenarios with lower input SISDR (Cohen et al., 2023).
@@ 4. Related Work
Commercial hearables. Conventional hearables use statistical signal processing for speech enhancement. However, deep neural networks have been demonstrated to achieve superior source separation, outperforming traditional methods by up to 9 dB (Subakan et al., 2021b; Luo and Mesgarani, 2019a; Chatterjee et al., 2022). Companies like Google (Park, 2024) have begun integrating AI accelerators, such as the Tensor A1 chip in Pixel earbuds, though specific details remain undisclosed, and its primary use appears to be for active noise cancellation. Similarly, in late 2024, Phonak (Pho, 2024) introduced AI-powered hearing aids, but their deep learning models appear to be heavily duty-cycled to conserve battery life. These commercial solutions remain proprietary, making it unclear whether they use programmable AI or fixed-function integrated circuits (ICs) (Tambe et al., 2023). This highlights the need for academic exploration into the design space of programmable, low-power speech AI for hearables and a better understanding of the challenges of running streaming deep learning models on hearables—a gap our work addresses.
@@
Low-latency speech processing. Applications have different latency constraints in that they either can wait to process an entire audio file or require much quicker responses. In augmented hearing, minimizing latency between input and processed output is crucial. However, this can degrade performance due to less available information for predicting the output (Wang et al., 2022b). Prior work has proposed architectures for low-latency speech tasks (Valin et al., 2020; Sato et al., 2024; Schröter et al., 2022; Veluri et al., 2024, 2023a), and some speech challenges focus on these low-latency models (Akeroyd et al., 2023). But, these are designed for devices with significantly higher clock frequencies, power budgets, and memory footprints than our target hardware and none of these neural networks meet the computational constraints of wireless hearables.
Deep learning with computational contraints. Common methods include quantization (Polino et al., 2018), pruning, and knowledge distillation (Nathoo et al., 2024a; Chen et al., 2018; Nathoo et al., 2024b). In the audio domain, these techniques have been applied to tasks such as keyword spotting (Zhang et al., 2018), speaker verification (Pavan et al., 2024), and sound event detection (Wu et al., 2024). Prior work has also explored these techniques for speech enhancement and denoising (Cohen et al., [n. d.]; Choi et al., 2021; Westhausen and Meyer, 2023; Nathoo et al., 2024a; Chen et al., 2018). However, these models are not designed to operate in real-time on hearable hardware. Moreover, they do not process time and frequency components as individual sequences—an essential feature of state-of-the-art enhancement models. (Cohen et al., 2024) proposes a quantized network, but it is neither causal nor real-time.
@@ 5. Limitations and Discussion
Recent research has explored techniques with wired headsets like semantic hearing (Veluri et al., 2023b), sound bubbles (Chen et al., 2024a), and target speech hearing (Veluri et al., 2024). While we do not demonstrate the feasibility of implementing these capabilities on wireless hearables, it is worth noting that some of these systems (Chen et al., 2024a; Veluri et al., 2024) rely on the dual-path model architecture that we optimize in this paper. This suggests a potential path for expanding to these capabilities in the near future.
@@
Future work could also integrate additional sensors, such as PPG and temperature sensors, to support non-speech applications like physiological sensing (Montanari et al., 2024). Although our hardware supports multiple microphones per device, our neural networks use only one. Exploring multi-microphone processing on a single hearable and across both ears could enhance performance and warrants further investigation.
@@
Finally, our paper focuses on behind-the-ear hearing aid form factor devices. The next step in this research would be integrating these speech AI models into earbud-form-factor devices, which is likely feasible due to the compact size of the target low-power AI accelerator.

### v2

@@ Abstract.
Enhanced hearing, quantization-aware training, hearing aids
$c$s: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learningccs: Computer systems organization Embedded and cyber-physical systemsccs: Human-centered computing Ubiquitous and mobile devicesjournalyear: 2025copyright: ccconference: The 31st Annual International Conference on Mobile Computing and Networking; November 4–8, 2025; Hong Kong, Chinabooktitle: The 31st Annual International Conference on Mobile Computing and Networking (ACM MOBICOM ’25), November 4–8, 2025, Hong Kong, Chinadoi: 10.1145/3680207.3765251isbn: 979-8-4007-1129-9/2025/11
@@ 1. Introduction
  • Tiny, low-power AI accelerator hardware (syn, 2025; gre, 2023) is inherently far more constrained than GPUs and general-purpose embedded CPUs like the Raspberry Pi. On the other hand, effective neural architectures for speech enhancement rely on components such as convolutions (Luo and Mesgarani, 2019b), LSTMs (Luo et al., 2020), transformers (Subakan et al., 2021), and state-space layers (Li and Chen, 2024). These models prioritize speech quality over real-time, on-device, or low-power constraints — indeed, transformers and state-space models exceed our target hardware’s runtime and memory limits. While efforts like TinyLSTM (Fedorov et al., 2020) and its successor TinyDenoiser (Rusci et al., 2022) aim for on-device deployment, their 25 ms algorithmic latency falls short of the low-latency requirements for enhanced hearing applications.
@@
  • Real-time on-device efficient neural network. Prior on-device models like TinyDenoiser (Rusci et al., 2022) are constrained by their limited complexity and inability to achieve effective performance with sub-10ms latencies. Instead, we begin with dual-path models that process audio in the time-frequency (TF) domain, treating time and frequency components as sequences using recurrent networks (Yang et al., 2024; Chao et al., 2024). These architectures achieve state-of-the-art performance by modeling both temporal and frequency relationships. However, they are computationally expensive, as running recurrent networks across each frequency prevents real-time operation. Thus, they cannot run in real time on our hardware. To address this, our proposed architecture compresses frequencies using a linear layer and replaces LSTMs with Gated Recurrent Units (GRUs). These architectural changes and others listed in §2.2.2 achieve real-time performance by reducing runtime while outperforming TinyDenoiser for the denoising task.
  • Hardware-software co-design. Floating-point speech denoising models cannot achieve real-time operation on our target accelerator hardware. To address this, we explore various quantized model configurations and use quantization-aware training (QAT) to simulate non-linear quantization errors and fine-tune the network to mitigate error propagation across layers. This narrows the performance gap between our mixed-precision and fully floating-point speech models.
@@ 2.1.1. Stacked printed circuits boards
  • PERIPH: A 4-layer rigid board contains peripheral devices, including an AP Memory APS256XXN 256 Mbit RAM chip and an SSM6515 ultra-low-power I2S DAC and amplifier. The AI accelerator can control power to the RAM chip using a load switch (FPF1204UCX) when the RAM is not needed. Additionally, this board includes a NOR flash (Macronix MX25UW12845G) and an IMU (BMI323), which are reserved for future research. A right-angled 1 mm pitch connector is used to connect a Phonak Audeo Marvel M Receiver-in-canal (RIC), allowing audio playback into the ear canal. The board connects to the MIC circuit using Molex 5050661022.
@@ 2.2. Efficient Streaming Neural Network
As shown in Fig. 3A, end-to-end latency is defined as the time for a single audio sample to travel from the microphone input buffer, through the speech enhancement network, and into the speaker output buffer. This latency consists of two components: algorithmic latency, introduced by factors such as chunk size, lookahead, and overlap-add; and hardware latency, which accounts for the computation and I/O time required to process each chunk (Wang et al., 2022).
@@ 2.2.2. Neural Network Architecture
Dual-window approach for low-algorithmic latency. As shown in Fig. 3A, to achieve low algorithmic latency, we process the input audio $x_{i}$ in small chunks of size $L_{C}$ (ms), with an additional lookahead of $L_{F}$ (ms) and a lookback of $L_{B}$ (ms); we set $L_{C}=6ms$, $L_{B}=6ms$ and $L_{F}=4ms$. Lookahead and lookback padding improve frequency resolution, which is important for enhancement performance (Wang et al., 2022). However, increasing the lookahead and lookback padding also introduce higher algorithmic latency.
@@
To eliminate the algorithmic latency introduced by lookback padding, we adopt a dual-window-size approach (Wang et al., 2022; Mauler and Martin, 2007), where the synthesis window is shorter than the analysis window, as shown in Fig. 3D. By discarding the lookback samples after the inverse DFT and before applying the synthesis window, the overlap-add process for the current chunk no longer depends on the next chunk’s output. We use a rectangular window for analysis and use a specific function for the synthesis window $s$, described below, valid for $L_{C}>L_{F}$, to achieve perfect signal reconstruction (Griffin and Lim, 1984):
@@
Cache state management. When processing a stream of continuous chunks, many values can be reused to avoid redundant computations. To achieve this, we maintain and update a cache state $h_{i}$ at every inference step. As shown in Fig. 3C, our cache state $h_{i}$ consists of four components: 1) Since the kernel of the first 2D causal convolution layer requires prior chunks, we can avoid recomputing the STFT frames of these prior chunks by caching and reusing them, as illustrated in the figure. 2) We store the output of the sequence of dual-path modules from previous chunks, allowing us to compute the causal 2D deconvolution more efficiently. 3) Since computing the ISTFT for the current chunk depends on previous chunks, we maintain a buffer for the intermediate outputs of the 2D deconvolution layer. 4) We store the hidden and cell states of the temporal unidirectional LSTM for each dual-path module. This allows us to fully leverage the long-term receptive field of the recurrent network.
@@ 2.2.3. Training for Real-world Generalization
In practice, each device runs its own network, processing audio independently. Thus, we train two separate networks — one for the left channel and one for the right.
Fine-tuning with motion and data augmentation. In the previous stage, the sound source and wearer’s head were assumed to be static. However, real-world scenarios involve motion. To address this, we fine-tune the model with time-varying $\theta$ and $\phi$, following motion simulation from  (Veluri et al., 2024).
Specifically, the source’s position updates every 25 ms, with a 2.5% probability of triggering a random motion event. When triggered, the angular velocity in the azimuthal and polar directions is sampled from $\pi/6,\pi/2$] rad/s, and the source moves at this velocity for a random duration between 0.1 and 1 s, during which no other motion event is triggered. The Steam Audio SDK (SDK, 2023) is used for motion trajectory simulation, creating a mix of stationary and moving segments within the same audio clip. Since BRIR datasets contain recordings at discrete positions, we approximate intermediate positions using nearest-neighbor BRIR selection based on the current $\theta(t),\phi(t)]$.
We also augment, with a 30 % probability, the mixture signal with white, pink, and brown noise, mimicking real-world noise sources like microphone thermal noise and HVAC systems. White noise is generated with a standard deviation sampled from [0,0.002), while pink and brown noise are created using the Python colorednoise library and scaled by a random factor from [0,0.05]. Additionally, also with a 30 % probability, we randomly augment the speed of the target speech to 80–120 % of its original value. We fine-tune the model for 100 additional epochs, incorporating motion simulation and noise augmentation. Instead of SNR loss, we use a multi-resolution spectrogram loss (Yamamoto et al., 2020; Steinmetz and Reiss, 2020). Optimization uses AdamW with a gradient clipping of 0.1, and a ReduceLROnPlateau scheduler (patience = 5, factor = 0.5) starting at 1e-3 learning rate.
@@ 2.3. Hardware-Software Co-Design
Efficient DSP and neural network inference rely on GAP9’s compute cluster — 9 RISC-V cores and the NE16 accelerator. The NE16 is optimized for streamed multiply-accumulate operations but only supports up to 8-bit quantized weights and 8-/16-bit activations. The RISC-V cores handle both integer and floating-point operations but are less efficient for neural networks. Thus, the network must be carefully designed by identifying the parts of the network that can be efficiently run on the NE16 with quantization without degrading the output speech quality.
@@
where $Q(r)$ is the fixed point representation of $r$ in quantization space. $\lfloor\cdot\rceil$ denotes the rounding of the input to the nearest integer value.
@@
A key aspect of uniform quantization is choosing the scaling factor $S$. It defines how a given range $\alpha,\beta]$ of real values $r$ is divided into discrete partitions (Gholami et al., 2022): $S=\frac{\beta-\alpha}{2^{b}-1}$. In our implementation, we use a Min-Max Moving Average Observer to determine $\alpha,\beta]$.
In asymmetric quantization, the range $\alpha,\beta]$ is not necessarily symmetric with respect to the origin, i.e., $\alpha\neq\beta$. In symmetric quantization, in contrast, a symmetric clipping range ($\alpha=\beta$) is used and its scaling factor $S$ is computed as, $S=\frac{\text{max}(-\alpha,\beta)}{2^{b-1}-1}$ (Gholami et al., 2022), with the zero-point set to $Z=0$.
@@ 2.3.3. Mixed-Precision Configuration
One method for quantizing a trained floating-point neural network is Post-Training Quantization (PTQ). This involves: 1) calibrating the quantization range using a representative dataset, and 2) quantizing all tensors in the network to a lower-precision format.
@@ 2.3.4. Quantization-Aware Training for Better Performance
Fine-tuning: The mixed-precision model is fine-tuned for 30 epochs using the FQSE framework (Cohen et al., [n. d.]).
@@ 3.1. Benchmark results
Applying mixed-precision quantization helps recover 0.36 dB of SISDRi for TinyDenoiser and 2.6 dB for our model. However, the performance loss compared to the floating-point model remains non-negligible. After QAT, TinyDenoiser achieves an SISDRi of 5.97 dB, while our model reaches 8.19 dB, demonstrating the effectiveness of QAT, especially for deeper and more complex networks. Our mixed-precision model with QAT achieves the best performance within hardware constraints. Considering runtime, memory, and power consumption, our model balances performance and memory efficiency while maintaining real-time requirements (<6ms<6~ms) and power constraints (71.6mW71.6~mW), trading some runtime and power efficiency for improved speech enhancement.
@@
In Fig. 4, we plot the output SISDR as a function of input SISDR for both the floating-point model and the mixed-precision QAT model. The results show that QAT reduces the performance gap between our floating-point and quantized models across all input noise levels. An interesting observation is that the performance gap increases with higher input SISDR. As input SISDR increases, the output speech quality becomes clearer, making quantization noise more dominant and its effects more pronounced compared to scenarios with lower input SISDR (Cohen et al., [n. d.]).
@@ 4. Related Work
Commercial hearables. Conventional hearables use statistical signal processing for speech enhancement. However, deep neural networks have been demonstrated to achieve superior source separation, outperforming traditional methods by up to 9 dB (Luo and Mesgarani, 2019a; Chatterjee et al., 2022). Companies like Google (Park, 2024) have begun integrating AI accelerators, such as the Tensor A1 chip in Pixel earbuds, though specific technical details of applications remain undisclosed. Similarly, in late 2024, Phonak (Pho, 2024) introduced AI-powered hearing aids. These commercial solutions remain proprietary, often provide limited technical detail, making fair and in-depth comparisons challenging. This highlights the need for academic exploration into the design space of programmable, low-power speech AI for hearables and a better understanding of the challenges of running streaming deep learning models on hearables—a gap our work addresses.
@@
Low-latency speech processing. Applications have different latency constraints in that they either can wait to process an entire audio file or require much quicker responses. In augmented hearing, minimizing latency between input and processed output is crucial. However, this can degrade performance due to less available information for predicting the output (Wang et al., 2022). Prior work has proposed architectures for low-latency speech tasks (Valin et al., 2020; Sato et al., 2024; Schröter et al., 2022; Veluri et al., 2024, 2023a), and some speech challenges focus on these low-latency models (Akeroyd et al., 2023). But, these are designed for devices with significantly higher clock frequencies, power budgets, and memory footprints than our target hardware and none of these neural networks meet the computational constraints of wireless hearables.
Deep learning with computational contraints. Common methods include quantization (Polino et al., 2018), pruning, and knowledge distillation (Nathoo et al., 2024). In the audio domain, these techniques have been applied to tasks such as keyword spotting (Zhang et al., 2018), speaker verification (Pavan et al., 2024), and sound event detection (Wu et al., 2024). Prior work has also explored these techniques for speech enhancement and denoising (Cohen et al., [n. d.]; Choi et al., 2021; Westhausen and Meyer, 2023; Nathoo et al., 2024; Chen et al., 2018). However, these models are not designed to operate in real-time on hearable hardware. Moreover, they do not process time and frequency components as individual sequences—an essential feature of state-of-the-art enhancement models. (Cohen et al., 2024) proposes a quantized network, but it is neither causal nor real-time.
@@ 5. Limitations and Discussion
Recent research has explored techniques with wired headsets like semantic hearing (Veluri et al., 2023b), sound bubbles (Chen et al., 2024a), and target speech hearing (Veluri et al., 2024). While we do not demonstrate the feasibility of implementing these capabilities on wireless hearables, it is worth noting that some of these systems (Chen et al., 2024a; Veluri et al., 2024) rely on the dual-path model architecture that we optimize in this paper. Further, recent models like TF-MLPNet (Itani et al., 2025) achieve promising results for speech separation using tiny quantized models. This suggests a potential path for expanding to these capabilities in the near future.
@@
Future work could also integrate additional sensors, such as PPG and temperature sensors, to support non-speech applications like physiological sensing (Montanari et al., 2024). Although our hardware has multiple microphones per device, our neural networks use only one. Exploring multi-microphone processing on a single hearable and across both ears could enhance performance and warrants further investigation.
@@
Our paper focuses on behind-the-ear hearing aid form factor devices. The next step in this research would be integrating these speech AI models into earbud-form-factor devices, which is likely feasible due to the compact size of the target low-power AI accelerator.
Finally, the presence of a real-time AI accelerator with sufficient I/O support not only enables enhanced hearing capabilities but also paves the way for a broad spectrum of augmented intelligence applications. These may include situational awareness, personal voice assistants, and cognitive support tools that extend beyond hearing enhancement. Thus, platforms such as ours that have embedded AI accelerators can enable broader research and development in edge-AI-powered intelligence augmentation.